3. A decision maker considers that the demand is affected by the season since it is highest during the winter. He aims at making a forecast for each season of the next year. Total annual demand is 4000 customers in the next year. Based on the information given below, what is the forecast for each season of the next year? Visitors Season Fall Winter Spring Summer Year 1 200 1400 520 720 Year 2 230 1600 580 831
Answer #1

Steps:

  • First, compute seasonal averages
    • Fall: (200+230)/2 = 215
    • Winter: (1400+1600)/2 = 1500
    • Spring: (520+580)/2 = 550
    • Summer: (720+831)/2 = 775.5
  • From these seasonal averages, find the seasonal indices.
    • Overall average = (200+1400+520+720+230+1600+580+831)/8 = 760.125
    • Seasonal index for Fall: 215 / 760.125 = 0.283
    • Seasonal index for Winter: 1500 / 760.125 = 1.973
    • Seasonal index for Spring: 550 / 760.125 = 0.724
    • Seasonal index for Summer: 775.5 / 760.125 = 1.020
  • Find the average season-wise forecast without any effect of seasonality. In other words, divide the yearly forecast into 4 equal forecasts (1,000 for each season).
  • Multiply the above flat forecast with the seasonal indices to get the adjusted forecast
Seasonal indices Average forecast (without seasonality) Adjusted seasonal forecast
0.283 1000 283
1.973 1000 1973
0.724 1000 724
1.020 1000 1020
Total 4000 4000

In Excel, we can implement the entire process as:

A B Seasonal Average forecast (without seas Adjusted seasonal forecast Year Season Visitors Seasonal indices average 31 F200D3+D7)/2 E3/AVERAGE SES3:SES6) 4 Winter 1400 D4+D8)/2 E4/AVERAGE (SES3:SES6) Spring 520 =(D5+D9/2 -ES AVERAGE(SE$3:$E$6) Summer 720D6+D10)/2 E6/AVERAGE (SES3:SES6) 2 Fall 230 Winter 1600 Spring 580 Summer 831 -F4 10 -F6 -F7 -F8 3 Fall -4000/4 -4000/4 -4000/4 -4000/4 SUM(G11G14) SUMHi1H14) -G11*F11 -G12*F12 -G13*F13 -G14*F14 Winter 12 13 14 15 Spring Summer -F10 Total